# Limit Theorems For Non-Hermitian Ensembles

### v3

@@ A.1. Proof of Theorem 6
The function $f_{N}(x)$ is chosen with the use of the definition of a class of limiting distributions exposed in the reference [12]. For a convenient choice of the function $f_{N}(x)$, the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for maxima.
@@
And, for any $\alpha>0$ and fixed $|x|\ll N$
$\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
$=\frac{1}{2}\frac{\log{N}}{\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
$=\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
$\sim\frac{1}{2}\frac{\log{N}}{\left[\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N\right]}$
$\sim\frac{\log{N}}{\log((1+\alpha)N/(2\pi))}$
$=\frac{1}{1+\frac{\log\left(\frac{1+\alpha}{2\pi}\right)}{\log N}}$
@@
Finally, let $R_{N}$ denote the scaled spectral radius.
@@
This limiting distribution is the standard Gumbel distribution of maxima. The corresponding cumulative density function is the function $F_{Gumbel}^{(max)}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is a random variable approximated as
@@
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$ with $\gamma_{\alpha N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$. The random variable $Z$ denote a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(maximum)-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ is Gumbel-distributed. This completes the proof of Theorem 6.
@@ A.2. Proof of Theorem 7
It is demonstrated in Section 2 that the scaled minimum radius $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ of eigenvalues for matrices from the complex Ginibre ensemble, close to zero, follows a Rayleigh distribution. The complex Ginibre ensemble does correspond to the complex induced Ginibre ensemble when the rectangularity index $L$ is equal to zero. Similarly stated in Section 3, Theorem 7 is as follows.
Theorem (Limiting distribution of the scaled minimum modulus $r_{N}$ for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Then,
where $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum modulus $r_{N}$ is approximated with the following
$r_{N}\simeq 1-T_{\alpha,N}+\xi_{\alpha,N}$ (44)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
Proof.
@@
$\approx\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{1}{2}\left(1+\operatorname{erf}\left(-\frac{\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)}{\sqrt{2}}\right)\right)\right]$
@@
$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
@@
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
As in the reference [20], it is assumed that $k=o(\alpha N)$. The factor $\sqrt{\frac{\alpha N}{\alpha N+k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$. For large values of $k$, this factor is rather large and the probability
$P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$ is approximatively equal to one. This means that its logarithm is zero.
@@
The variable $\bar{Z}=\frac{1}{\alpha N+k}\sum_{j=1}^{\alpha N+k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is
$\frac{\sigma}{\sqrt{\alpha N+k}}=\frac{1}{\sqrt{\alpha N+k}}$ with $\sigma=1$.
@@
$=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+K_{N}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N+k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)+O\left(\frac{1}{\alpha N+k}\right)$
$+O\left(K_{N}(\alpha N+k)^{-3/2}\right)+O\left(e^{-\frac{1}{2}K_{N}^{2}}\right)$
where, similarly stated in [20], for $x$ restricted as in $|x|<Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
@@
where $E_{N}$ is the error term
@@
The term $\frac{1}{\sqrt{\alpha N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left(\alpha N\right)}\right)$. This implies the following result
@@
The integral $\int_{-\infty}^{-t}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$.
Thus, the logarithm of the limiting survival distribution function $1-F_{\infty}(x)$ is
@@
and in the limit as $N$ goes to infinity
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$
@@
which implies that
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\left(\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right)\right]$
$=\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
$=\exp(-\exp(x))$
@@
and corresponds to distribution function of the standard Gumbel distribution for minima.
@@
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$ with $\gamma_{\alpha N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$. The random variable $Z$ denotes a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(minimum)-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ is Gumbel-distributed. This completes the proof of Theorem 7.
@@ A.3. Proof of Theorem 10
The independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble is established from Theorem 10 presented here again for more convenience.
Theorem (Independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$. The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity. More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
Proof.
The independence of the scaled spectral radius and the scaled minimum modulus is derived from their joint cumulative distribution function
@@
Considering the scaling $\sqrt{\alpha N}$, the inner radius $r_{in}=1$ and the outer radius $r_{out}=\sqrt{\frac{1+\alpha}{\alpha}}=\gamma$. Let $r=1-\frac{f_{N}(y)}{\sqrt{\alpha N}}$ and $R=\sqrt{\frac{1+\alpha}{\alpha}}+\frac{f_{N}(x)}{\sqrt{\alpha N}}=\rho+\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ and $f_{N}(y)$ are $o\left(\alpha N\right)$.
@@
$=\prod_{k=1}^{N}\left[P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\right]$
where $C=\rho^{2}-1>0$. The functions $\phi_{N}^{(r)}(y)=2f_{N}(y)\left(1-\frac{f_{N}(y)}{2\sqrt{\alpha N}}\right)$ and
$\phi_{N}^{(R)}(x)=2\rho f_{N}(x)\left(1+\frac{f_{N}(x)}{2\rho\sqrt{\alpha N}}\right)$.
Furthermore, with proportional rectangularity index $L=\alpha N$, $\alpha>0$,
$\prod_{k=1}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\times$
$\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
$=\prod_{k=1}^{N}p_{k}\prod_{k=1}^{N}h_{k}$
@@
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
and
$h_{k}=\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
The partial product $\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [20]), such as
$\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\leq\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
The following is derived using the same arguments as stated in [20] and using the Markov inequality and the definition of the quantile of the exponential distribution.
The function $\phi_{N}^{(R)}(x)$ is a positive and increasing function.
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
The random variables $Z_{j}$, $i\in\{1,\cdots,\alpha N+k\}$ are independent and identically distributed.
Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{(\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)+\alpha N\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{\alpha N+k}Z_{j}>\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)\right]$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left[1-e^{-\eta\sqrt{\alpha N}\phi_{N}(x)}E\left[e^{\eta Z_{1}}\right]^{\alpha N+k}\right]$

### v4

@@ A.1. Proof of Theorem 6
Then, a convenient choice of the function $f_{N}(x)$ is performed, such that the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for maxima. More precisely, choosing
@@
Also,
$\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}\\ =\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
And, for any $\alpha>0$ and fixed $|x|\ll N$,
$\lim_{N\rightarrow+\infty}\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
$=\lim_{N\rightarrow+\infty}\frac{1}{2}\frac{\log{N}}{\left[\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N\right]}$
$=\lim_{N\rightarrow+\infty}\frac{\log{N}}{\log((1+\alpha)N/(2\pi))}$
$=\lim_{N\rightarrow+\infty}\frac{1}{1+\frac{\log\left(\frac{1+\alpha}{2\pi}\right)}{\log N}}$
@@
Finally,
@@
This limiting distribution is the standard Gumbel distribution of maxima.
The scaled spectral radius $R_{N}$ is a random variable approximated by
@@
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$ with $\gamma_{\alpha N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$. The random variable $Z$ denote a random variable following a standard Exponential distribution. The random variable $X=-\log(Z)$ is a standard Gumbel(maximum)-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ is Gumbel-distributed.
This completes the proof of Theorem 6.
@@ A.2. Proof of Theorem 7
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\frac{\sqrt{\alpha N}}{\sqrt{\alpha N+k}}\phi_{N}(x)-\frac{k}{\sqrt{\alpha N+k}}\right)$
$>\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\frac{k}{\sqrt{\alpha N+k}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{1}{2}\left(1+\operatorname{erf}\left(-\frac{\left(\frac{k}{\sqrt{\alpha N+k}}\right)}{\sqrt{2}}\right)\right)\right]$
@@
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$
and, for fixed $k\in\left[0,\sqrt{2\alpha N\log(N)}\right]$, the factor $\sqrt{\frac{\alpha N}{\alpha N+k}}$ goes to one as $N$ goes to infinity.
@@
The variable $\bar{Z}=\frac{1}{\alpha N+k}\sum_{j=1}^{\alpha N+k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is equal to $1$ and its standard deviation is $\frac{\sigma}{\sqrt{\alpha N+k}}=\frac{1}{\sqrt{\alpha N+k}}$ with $\sigma=1$.
Applying the Edgeworth expansion, as in [22], the logarithm of the probability $p_{k}$ is then
@@
$=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+K_{N}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)+O\left(\frac{1}{\alpha N}\right)$
$+O\left(K_{N}(\alpha N)^{-3/2}\right)+O\left(e^{-\frac{1}{2}K_{N}^{2}}\right)$
where, similarly stated in [22], for $x$ restricted as in $|x|<H$ for some large positive $H$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
@@
where, with respect to the error terms $O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)$ and $O\left(\frac{1}{\alpha N}\right)$ of the Edgeworth expansion of the logarithm of the probability $p_{k}$,
@@
The term $\frac{1}{\sqrt{\alpha N}}$ appearing in the lower limit of integration is negligible. Extending the upper limit to $\infty$, the function $\phi_{N}(x)=o\left(\sqrt{\log\left(\alpha N\right)}\right)$. This implies the following result
@@
Thus, as the function $f_{N}$ goes to infinity, the logarithm of the limiting survival distribution function $1-F_{\infty}(x)$ is
@@
and in the limit as $N$ goes to infinity,
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}+O\left(\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{5}}dt\right)$
@@
This implies that
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\left(\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right)\right]=\exp(-\exp(x))$
@@
and corresponds to the distribution function of the standard Gumbel distribution for minima.
@@
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$ with $\gamma_{\alpha N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$. The random variable $Z$ denotes a random variable following a standard Exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ is Gumbel-distributed.
This completes the proof of Theorem 7.
@@ A.3. Proof of Theorem 10
The independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble is established as follows. The independence of the scaled spectral radius and the scaled minimum modulus is derived from their joint cumulative distribution function
@@
Considering the scaling $\sqrt{\alpha N}$, the inner radius $r_{in}=1$ and the outer radius $r_{out}=\sqrt{\frac{1+\alpha}{\alpha}}=\rho$. Let $r=1-\frac{f_{N}(y)}{\sqrt{\alpha N}}$ and $R=\sqrt{\frac{1+\alpha}{\alpha}}+\frac{f_{N}(x)}{\sqrt{\alpha N}}=\rho+\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ and $f_{N}(y)$ are $o\left(\sqrt{\alpha N}\right)$.
@@
$=\prod_{k=1}^{N}\left[P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\right]$
where $C_{\alpha,N}=\sqrt{\alpha N}\left(\rho^{2}-1\right)>0$. The functions $\phi_{N}^{(r)}(y)=2f_{N}(y)\left(1-\frac{f_{N}(y)}{2\sqrt{\alpha N}}\right)$ and
$\phi_{N}^{(R)}(x)=2\rho f_{N}(x)\left(1-\frac{f_{N}(x)}{2\rho\sqrt{\alpha N}}\right)$.
Furthermore, with proportional rectangularity index $L=\alpha N$, $\alpha>0$, the partial product
$\prod_{k=1}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]), such as
$\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\leq P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
@@
$p_{k}=P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$